OpenStudy (anonymous):

URRGGEENNTT!!! :) Solve the following system using the substitution method. 16x + 3y = 0 8x + y = 1

OpenStudy (anonymous):

16x + 3y = 0 8x + y = 1

OpenStudy (anonymous):

y=-8x+1

OpenStudy (anonymous):

plug that in first equation wherever you see y

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

(-3/8, 2)....(3/8, -2)....(-3/8, -2)....(3/8, 2) those are the answers. i cant figure out which one it is.

OpenStudy (anonymous):

16x + 3y = 0 plug that in first equation wherever you see y y=-8x+1 16x+3(-8x+1)=0 solve for x x=3/8

OpenStudy (anonymous):

y=-8(3/8)+1 y=-3+1=-2

OpenStudy (anonymous):

(3/8,-2)